### v2

# Neil A. Chriss

Abstract
This is the third paper in a series concerning the game-theoretic aspects of position-building while in competition. The first paper [Chr24a] set forth foundations and laid out the essential goal, which is to minimize implementation costs in light of how other traders are likely to trade. The majority of results in [Chr24a] center on the two traders in competition and equilibrium results are presented. The second paper, [Chr24b], introduces computational methods based on Fourier Series which allows the introduction of a broad range of constraints into the optimal strategies derived. The current paper returns to the unconstrained case and provides a complete solution to finding equilibrium strategies in competition and handles completely arbitrary situations. As a result we present a detailed analysis of the value (or not) of trade centralization and we show that firms who naively centralize trades do not generally benefit and sometimes, in fact, lose. On the other hand, firms that strategically centralize their trades generally will be able to benefit.
Keywords Trading, position-building, implementation cost, alpha-decay, game theory, equilibrium, price of anarchy, trading strategies, order-flow, front-running, centralized trading, naive centralization, strategic centralization
1 Introduction
Game theory is a framework for analyzing competitive interactions among players whose fortunes depend on the actions of all participants involved. One of its central concepts is Nash equilibrium, the set of strategies that are simultaneously optimal with respect all of the others. The focus of this paper is to give a complete analysis of Nash equilibrium when the strategic interactions are two or more traders simultaneously purchasing the stock of a particular company over a fixed period of time. This continues the work in [Chr24a] which introduced the concept of trading in competition, defined standard terminology and found Nash equilibria for two-trader general and multi-trader symmetric equilibria, The current paper completes the work of [Chr24a] by providing a complete characterization of multi-trader equilibria. As an application we demonstrate the ways that firms who employ multiple traders can centralize their trades and disguise their order flow to lower trading costs. The rest of this introduction provides a high-level overview of the main results of this paper.
1.1 Competition, implementation cost and equilibrium
What drives competition in trading and what are the traders competing for? We examine the situation in which two or more traders simultaneously believe that there is a catalyst that at some nearby future time will cause the price of a stock to rise to a new price called the target price [^1] This is a positive catalyst. There are obviously negative catalysts as well and the work here works identically for those. We will sometimes refer to these traders as the informed traders or just "the traders" when there is no risk of confusion. The difference between the current price and the target price represents a theoretical profit that is the motivation for traders to purchase the stock. Traders individually choose their target quantities to purchase and they all would like to capture as much of the theoretical profit as possible. Though these traders represent a minority of all traders who might trade in the stock, we model them as the only traders acting on the information related to the catalyst. Therefore their collective buying represents the marginal demand for the stock and puts upward pressure on its price. We presume all traders recognize this and therefore select the sizing and timing of their trades to minimize the price they part for the stock. Standing in the way are two forms of market impact, those effects that cause the price of the stock to move as a result of their collective trading.
The first market impact effect is the persistent rise in price due to the collective and persistent demand for the stock from the informed traders. This persistent buying "leaks" into the market and causes other participants to re-rate the stock’s value upward. This is often referred to as permanent market impact or alpha-decay where in this context the former means permanent only until the catalyst occurs and the latter refers to the degradation of theoretical profit due to the trading. For a given trader, the cost of acquiring their target position is measured as the difference between the price per share of the stock immediately before trading begins and the average price paid is called the implementation cost of the trade.
Each trader wishes to minimize their implementation cost by choosing the timing and sizing of their trades, the totality of is called a trading strategy. Each trader must select their strategy taking into account what their adversaries strategies are likely to be. One way to imagine this is that each trader is shifting portions of their target quantity backward and forward in time in order to balance two opposite considerations. Shifting trades earlier in time lowers later purchase prices by avoiding alpha-decay, but this comes at the expense of higher temporary impact. Determining the optimal mix depends on all the strategies but because traders are not coordinating wit h one another they need to select their strategies according to general principles. A Nash equilibrium (see [Nas50]) is the key concept in this regard.
1.2 The market strategy and equilibrium strategies
A Nash equilibrium is a set of strategies that simultaneously minimize every trader’s implementation cost. By definition, in such an equilibrium if any trader modifies their strategy they will increase their implementation cost and therefore an equilibrium is theoretically stable in the sense that each trader has no incentive to modify their strategy. Section 5 proves that in any multi-trader competition for a single stock there is a unique set of equilibrium strategies that have closed-form descriptions for both the schedule of trades as well as the implementation costs. We describe this next.
In [Chr24a] we did not provide any general statements about equilibrium. It turns out that in order to find these requires a new ingredient. The market strategy, discussed in Section 3, is simply the aggregate of all traders strategies. Though a strategy usually means that the sizing and timing of trades is being controlled by one individual, we nevertheless treat the aggregate trading as if it is a strategy in the conventional sense. Treating the market strategy as an actual strategy allows us to study its properties when its constituent strategies are in equilibrium. A key result is that in an equilibrium the market strategy is an eager strategy, that is, it acquires stock more aggressively than a straight-line strategy that maximally spreads its trades out over time (see Section 1 for more details).
With this in hand a number of new results become available. First, total implementation cost of the market strategy, or equivalently, the sum of the cost to each trader is independent of how distributed and only depends on the number of traders and the temporary impact and alpha-decay parameters. This arises because there is a highly constrained relationship between the the individual strategies and the market strategy. Not only does this allow us to provide equilibrium strategies in closed-form (see Section 5) but also provide formulas for the implementation cost of each strategy, e.g., see Proposition 5.2. From this we immediately see another surprising fact: each traders implementation cost only depends on their target quantity as a fraction of all the other traders.
1.3 Centralized trading and disguising order flow
Modern firms such as multi-strategy hedge funds and mutual funds often have multiple traders who are trading in competition for the same stock despite working for the same firm. One can model this situation without regard to the firm itself, but we can also ask whether the firm should consider whether it can, in some, way consolidate trades to its advantage. In Section 6 we use the equilibrium strategies and cost formulas we derive to analyze the potential benefits of trade centralization from a game-theoretic standpoint, taking into account that there are still traders outside of the firm competing for the same stock. We find that if a firm opts to naively centralize and trade their order flow as a single, monolithic trade it may very well be worse off centralizing. Alternatively, a firm may strategically centralize its trades by disguising its true number of traders and in this case it has material opportunities to lower its implementation costs. In the former case, the reason firms will find it challenging to benefit from naive centralization is because in this case its orders will become large and vulnerable to front-running as described in Section 1. Naive centralization does provide benefits overall, but ironically they accrue to traders who are not in the firm. However, if a firm strategically centralizes, that is, if they take their consolidated order flow and work to mask the true number of traders and optimize, they there can be substantial improvements to implementation costs.
2 Detailed background information
In this section we review the key concepts used through this paper in detail, repeating and enriching the overview in the introduction. For the most part these concepts were introduced in [Chr24a], however several, including the market strategy and trading coalitions, are new.
2.1 Catalysts, urgency and competition
This paper concerns trading in competition so the first thing to settle is why there is competition and how does this shape demand for a stock. Our idealized model begins with the concept of a catalyst that will occur at a fixed future date that has the power to significantly revalue the company and cause the stock price to significantly change. This is a fairly prevalent situation in equity markets world-wide; hedge funds and similar market participants do research to find catalysts and then build positions to capture the profit available from the price move. There are several important features of this setup that drive the need to explore a game-theoretic analysis.
First, we assume there is a relatively small number of informed traders who are aware of the catalyst and want to build positions in order to profit from it. These traders are collectively setting the price for the stock because their aggregate demand is creates a persistent imbalance in demand for the stock, relative to the its supply. Because there is a fixed date on which the catalyst will occur, traders have to acquire their positions prior to the catalyst date in order to earn their desired profit. This creates urgency because under the implicit rules of this game, traders are competing with other trades for liquidity and have limited flexibility to modify how they trade in light of the demand imbalances they collectively create. Thus traders need a strategy for how they are going to build their position that takes into account the likely strategies of the other traders. Thus to discuss this further we will need pin down exactly what we mean by "trading strategy".
More formally, we say the number of informed traders (henceforth, just traders) is $n$ and identify each trader by $i=1,\dots,n$ and refer to the trader whose index is $i$ as the trader $i$. Each trader has a target quantity of stock they wish to acquire and we write $\lambda_{i}$ for the portion of the total quantity their target quantity represents. Each trader’s strategy is represented by $\lambda_{i}a_{i}$ where $a_{i}$ is a function that says how much cumulative quantity the trader will have acquired by time $t$. This is formally a trading strategy and we discuss this next.
2.2 Trading strategies
In order to studying trading strategies precisely we idealize the situation as follows. We assume there are $n$ traders who all start to trade at a fixed time $t=0$; they all commence trading at this time and measure their costs relative to the prevailing price at the time which is variously the reference price or arrival price. Prior to this time each trader each decides on their target quantity and we normalize all target quantities to 1. We write $\lambda_{i}$ for trader $i$’s target quantity as a fraction of the whole so that $\sum_{i=1}^{n}\lambda_{i}=1$. Thus target quantity represents the portion of the total quantity a trader seeks.
Each trader’s target quantity is regarded as a single order to buy the stock, and then the trader breaks that em parent order into subsidiary orders tied to it called child orders. A trading strategy is the precise scheduling of child orders so as to implement the parent order.
  • Trading strategies always begin at a time denoted $t=0$ and end at time $t=1$. Thus time is scaled, this makes some of the mathematics significantly more straightforward and sacrifices no generality;
  • A trading is called a position-building strategy if it starts with no stock (at $t=0$) and ends with a positive quantity of stock (at $t=1$);
  • We only consider position-building strategies and thus use the term trading strategy and position-building strategy analogously;
  • A position-building strategy whose target quantity is $1$ is called a unit strategy; and
  • When two or more traders are building a position in the same stock over the same interval we say they are building the positions in competition. This is the primary focus here and in [Chr24a].
  • When a trader is building a position in a stock, we will refer to this as taking up the trade toward acquiring the target quantity.
  • The position–building strategy is specified with two components: the target quantity $\lambda_{i}$ and the unit strategy $a_{i}$ where $i$ is the index of the trader.
Thus a trading strategy is a strategy for acquiring a target quantity of stock starting at a fixed date and ending with a desired quantity of stock prior to the catalyst date. In fact, it is a precise schedule of trades that says how much to trade at each point in time, which, naturally is a key limitation of the framework here and [Chr24a]. As set forth the strategies do not anticipate incorporating market information into the model as time evolves. Therefore the strategies obtained here are most useful in precisely those situations in which the basic contours of trading are not likely to change during the course of trading, which we argue is often the case.
2.3 A taxonomy of trading strategies
Since every trading strategy is represented by a function, it is possible to divide strategies into by the shape of their functions. Strategy shapes we discussed in [Chr24a] but for completeness we review the key points here. There are three categories of simple strategies, those whose second derivative is constant sign: risk-averse, risk-neutral and eager according to whether their second derivative is positive, zero or negative. Figure 1 gives a visual display of these:
Figure 1: The three simple strategy shapes when second derivative is constant.
A different categorization of strategies is called aggressive strategies. These are strategies acquire more rapidly than the market-neutral strategies and do so in order to front-run other strategies in order to, in effect, their market impact their competitors (see Section 2.4 for more on market impact). Several examples of aggressive strategies are shown in Figure 2 and these are quite common in equilibrium settings.
Figure 2: Aggressive strategies, only one of which is simple in that it never overbuys and therefore never sells, while the other two do.
In the simplest terms, front running occurs when a trader uses advance knowledge of an upcoming large transaction to exploit market movements to their advantage. In essence, the front runner anticipates the inevitable market reaction to a large order, when the rest of the market adjusts. From a game-theoretic perspective, front running can be seen as a strategy that seeks to exploit a trader’s knowledge of what other traders are likely to do in order to adjust their strategies accordingly.
2.4 Implementation cost and market impact
Here and throughout, we assume as usual that there is a small number of informed traders competing to buy a stock. We assume that each trader has selected a strategy and ask, what will it cost each trader to acquire their position? This cost has a specific meaning: it is the in excess of the reference price, that is, the price paid (per share) in excess of the prevailing price prior to the start of trading. At work is the assumption that the aggregate demand for liquidity caused by the traders outstrips supply, for example caused by their collective urgency, and the persistent supply-demand imbalance causes a gradual revaluation of the stock due to the presumed informativeness of the trades. In the whole, the price average execution cost of the acquisition is called the implementation cost and, of course, traders are expected to choose their strategies in such a way as to reduce their implementation cost.
We call the former component of market impact temporary impact and the latter as permanent impact or when we are referring to the erosion of theoretical profit, alpha-decay. Both temporary and permanent impact have been studied extensively from both theoretical points of view, including detailed analyses of order book dynamics and market microstructure. In the optimal execution and microstructure literature the persistent component.
Modeling market impact
We briefly explain how we model market impact in this paper. To start, at any given time trader $i$’s cumulative stock acquisition through time $t$ is given by its strategy, the function $\lambda_{i}a_{i}(t)$. Therefore their change in position between two nearby times $t_{1}$ and $t_{2}$ is $\lambda_{i}(a_{i}(t_{2})-a_{i}(t_{1}))$ which is the quantity of stock traded between those two times. Should we divide by $t_{2}-t_{1}$ we arrive at the rate of trading. We model the impact on the price of a stock due to immediacy, the temporary impact, as proportional to the rate of trading [^2] This is the standard linear temporary impact model studied in the optimal execution literature, see [AC97] and [AC01] for example..
As an example, if a trader sets units of time to one hour and trades at 5,000 units (e.g., dollars) per hour then we are are modeling the price impact as proportional to 5000 units per hour [^3] As we are using these models to select strategies the constant of proportionality is irrelevant. If we were using these models to accurately forecast price changes we would need to precisely calibrate the model for that purpose. Since the trader also is buying the 5,000 units, the total increased price paid is proportional to $000\times 5,$ 000 × 5 , 000 units. Thus the impact on implementation cost is quadratic. Finally, it is important to note that we will be modeling the aggregate impact that all of the traders trading in competition have on the market, and any one in particular.
We model permanent impact quite differently based on the cumulative quantity acquired. In this case, our model reflects the notion that as the informed traders collectively trade, the total quantity they acquire is proportional to the level of information the collectively "leak" into the market and the impact on price is proportional to that. We model this as permanent impact is proportional to the cumulative stock purchased by all of the traders through a given time. We write $\kappa$ for the constant of proportionality and call it the market impact parameter; $\kappa$ controls the relative contribution of permanent impact versus temporary impact.
The market strategy and market impact
We write $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ for the implied trading strategy due to the aggregate trading of all the competitive traders. We call this the market strategy (this is neither the market nor a strategy however it is elaborated on extensively later in this paper). Because the competitive traders represent the marginal supply and demand for the stock their collective trading is the key driver of market impact. We model the instantaneous cost of trading as proportional to the market strategy’s rate of trading and the permanent impact as proportional to the cumulative trading by the market strategy. Both of these model the change in stock price, the former relative to the prevailing market price and the latter relative to the reference price. Adding these up, says that the total cost at time $t$ is proportional to:
$C_{i,\lambda_{i}}(t)\equiv\dot{m}(t)\,\lambda_{i}\dot{a}_{i}(t)+\kappa\,\lambda_{i}\,m(t)\dot{a}_{i}(t)$ (2.1)
This value represents the cost relative to every trade occurring at the prevailing price without any permanent impact due to trading and only up to a constant multiple. Thus it is not a measure of the actual cost of trading at time $t$; however, since our aim is to find optimal, that is lowest cost, strategies, we are only interested in cost up to a constant.
2.5 Minimizing implementation cost
The instantaneous cost function in Eq. 2.1, $C_{i,\lambda_{i}}$ takes into account the trading of all traders and both the temporary impact and alpha-decay cost of trading. The implementation cost to trader $i$ is the integral from $t=0$ to $t=1$ of the cost function:
$\text{Cost}(a_{\lambda_{i}},m)=\int_{0}^{1}C_{i,\lambda_{i}}(t)\,\mathrm{dt}=\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\lambda_{i}a_{i}(t)\,\mathrm{dt}$ (2.2)
Equation (2.2) is the total implementation cost to trader $i$ of trading in competition with traders $i=1,\dots,n$ whose market strategy is $m=\sum\lambda_{i}a_{i}$. This is the juncture at which game theory enters the scene because each trader would like to minimize their implementation cost with respect to the others. Thus, each trader needs to make a choice based on what the others are likely to do.
A tool we will use repeatedly to find strategy that minimize implementation cost is the Euler-Lagrange equation. This states that a function $a_{i}:[0,1]\to\mathbf{R}$ that minimizes the above integral satisfies the following linear partial differential equation:
$\frac{\partial}{\partial t}\frac{\partial C_{i,\lambda_{i}}}{\partial\dot{a}_{i}}-\frac{\partial C_{i,\lambda_{i}}}{\partial a_{i}}=0$ (2.3)
Since our aim is to find Nash equilibria, our aim is therefore to find strategies $\lambda_{1}c_{1},\dots,\lambda_{n}c_{n}$ such that each trader’s strategy simultaneously minimizes their cost function, which clearly involves finding the simultaneous solution to Eq. 2.3 for all $i$.
2.6 Future directions
Finally, we give a brief discussion of various directions one might consider. This paper’s primary aim is to shed light on strategy decision making while trading in competition and we believe a game-theoretic framework is the right starting point. However, it is obvious that this framework has many shortcomings that provide a multitude of potential research directions. The good news is that the historical development of game theory itself provides a useful template for how one might evolve this research.
Game-theoretic directions
The $n$-trader equilibria developed in this paper are described formally as an $n$-player, simultaneous one-shot, perfect information, non-cooperative games. We describe each of these briefly here:
  • One-shot games: players make decisions simultaneously and only once, without the opportunity for future rounds or revisions. In these games, players must consider their strategies and payoffs based on their expectations of the other players’ choices.
  • Simultaneous games: players make their decisions at the same time without knowing the choices of the other players. In these games, each player must decide based on their beliefs or expectations about the others’ actions, as no player has information about others’ moves when making their own.
  • Perfect information games: all players have complete knowledge of the game including the strategies every other player will play. In these games, players make decisions sequentially, fully aware of all previous moves, allowing them to optimize their strategy with complete information about the current state of the game.
  • Non-cooperative games: players pursue their individual interests, often competing against one another, and the outcome depends on the strategic choices of all participants, leading to an equilibrium where no player can improve their payoff by unilaterally changing their strategy. The Nash equilibrium is a key concept in analyzing non-cooperative games.
Each item in this list marks a potential direction of research by going beyond our framework in a direction that tracks the development of game theory since Nash. For example, traders repeatedly trade in competition, so to what extent would modeling this as a repeated game impact equilibria? Also, in reality traders do not "make their moves" simultaneously, but rather different traders not only start their position-building at different times, but at times they go so far as to announce their ideas in conferences, public interview and private "idea dinners" with other investors. In addition, modeling this as a perfect information game is decidedly not a perfect information game. Traders do not know what the other traders’ strategies are or what moves have occurred up to a given time.
One game theoretic direction that seems particularly important to explore is Bayesian game theory, see [Har67]. Bayesian game theory extends classical game theory to situations with incomplete information, where players (i.e., traders) do not have complete information about the other players strategies which could include the form of the strategies themselves, target size or restrictions a trader might place on their strategy. Put differently, Bayesian game theory provides a framework for studying games where players are uncertain about other players strategies. In this setup, each player (read: trader) forms beliefs about the other players’ strategies based on a common prior distribution, updating their beliefs according to Bayes’ rule to maximize their expected utility (which in this case means minimizing their expected losses).
Market impact models
In addition to the game theoretic limitations, there is also the question which market impact models should one use to inform trading in competition? At this point it may seem like the only right answer is to use models that most accurately predict the impact of trading on prices, however a counterpoint to this is that since we are not, in fact, interested in predicting market impact or implementation cost itself, it is not wholly obvious that this is the right answer. One reason for this is because the details of market impact, especially alpha-decay will vary over time and must be estimated each time. Therefore, it is worth considering robust models that prioritize minimizing the cost of errors in parameters over the specificity of the models. Despite this, it is still worth examining the how other models might change the shape of strategies in equilibrium and importantly whether or not the nature of equilibrium itself is dependent on the market impact models used.
There is an enormous academic literature on market impact. A sample of its breadth includes market-microstructure theory (the mechanisms and processes by which securities are traded and how these factors influence price formation, liquidity, and transaction costs), empirical investigations of market impact functions, to the study of optimal execution that studies trading strategies, ordinarily in non-competitive situations. Of particular interest are non-linear market impact models and decay kernels that describe how the impact of trading in the near-term effects prices and how this impact dissipates over time. Many modern treatments of market impact involve price impacts that fit within a general model described as follows. We let $K$ be a function called a decay kernel that roughly speaking describes how the impact of trading dissipates over time, $\mathcal{I}$ be an impact function that describes the functional form of the impact of the rate of trading in a stock on its price and $a$ a trading strategy. Then the change in price is modeled as:
Cumulative $\textit{Cumulative impact}=\int_{0}^{t}K(t-s)\,\mathcal{I}(\dot{a}(s))\,\mathrm{ds}$ (2.4)
This model is sufficiently general as to capture a rich array of market impact models, including the ones used here and in [Chr24a] (as well as the original work on optimal execution in [AC97] and [AC01]). More broadly, the above model is general enough to describe as special cases some of the key models that have been studied empirically and theoretically including [Gat10a], [GS11], [GS13], [OW13] and [BFL09]. Also see [FGLM06], [ATHL05] and [ZTFL15] [^4] This literature is, in fact, much broader than we can do justice to here, but [Gat10b] and [Web23] are excellent references.
In general, it would be interesting to investigate how changing to different market impact models influences the nature of equilibrium. Does it change the uniqueness of equilibrium? A step in this direction is possible using the computational methods developed in [Chr24b] which develops methods using Fourier series that make make computing optimal strategies and equilibria more computational tractable in certain situations. This paper was developed to add real-world constraints to strategies in competition, these methods may also be applied to studying optimal trading in competition with other impact models, e.g., the model in [OW13].
Mathematical models
A key assumption in this paper is that the functions describing the trading strategies are twice differentiable. This makes it possible to study equilibrium using variational methods and fairly straightforward differential equations. It also allows for very straightforward approximation methods using Fourier series as in [Chr24b]. This convenience, however, comes at the price of not being able to include trading strategies that are not differentiable. On the one hand this might not be so bad because we can approximate non-differentiable functions arbitrarily closely with differentiable ones. On the other hand, when it comes to discontinuities, it is at least worth speculating whether allowing for discontinuities is necessary. Discontinuities in trading strategies represent block trades, including privately negotiated transactions executed outside of the open market, traded to minimize its impact on the stock’s price due to its size. It is natural to ask whether allowing for block trades and potentially different impact dynamics for those trades would yield different and potentially valuable insights regarding trading in competition.
2.7 Organization of this paper
Here we give a brief overview of the rest of this paper. In Section 3 we examine what happens when all traders’ strategies are consolidated into a single strategy. We do this so that later we can examine how the associated costs of a market minimized strategy compare with those in equilibrium. Next, in Section 4 we study the market strategy, the aggregation of all strategies treated as if it were a strategy, and we study its properties at equilibrium. In Section 5 we prove that equilibrium strategies exist and in Proposition 5.1 we provide closed form solutions for those strategies. This section also derives formulas for implementation costs in equilibrium. Finally, in Section 6 we examine trade centralization and study how and to what extent it can improve implementation costs for a firm that employs multiple traders. Finally, we relegate most of the proofs of purely mathematical propositions to a set of appendices at the end of this paper.
3 Whole market cost minimization
In this section we examine what happens when $n$ traders wish to buy the same stock but rather than trading in competition their trades are aggregated into a single trade and they are implicitly forced to trade cooperatively. Such might be the case, for example, if the traders work for the same hedge fund and all trader’s trades are sent to a central trading desk that behaves as a single monolithic trader. The punchline is that in such a case and in the absence of any aversion to market risk the centralized trades should be executed as a risk-neutral strategy.
We examine this here so that later we can analyze the price of anarchy referred to in the introduction. While the result is immediate if we simply treat the market strategy as a single strategy, we show here its precise derivation as the sum of the implementation costs of each strategy. Let $C$ be the total cost of trading (function) equal to the sum of the individual costs $C_{i,\lambda_{i}}$. In that each $C_{i,\lambda_{i}}$ is equal to $(\dot{m}+\kappa m)\lambda_{i}\dot{a}_{i}$ we immediately see
$C=\sum_{i=1}^{n}(\dot{m}+\kappa m)\lambda_{i}\dot{a}_{i}=\dot{m}^{2}+\kappa m\,\dot{m}$ (3.1)
This relation depends on the fact that the market impact is linear. Namely for two general strategies $x$ and $y$ the cost of trading $x$ and $y$ simultaneously is equal to the sum costs of trading each. With that said we apply the Euler-Lagrange equation to find the strategy that minimizes implementation cost.
$\frac{\partial}{\partial t}\,\frac{\partial C}{\partial\dot{m}}=2\ddot{m}+\kappa\dot{m},\qquad\frac{\partial C}{\partial m}=\kappa\,\dot{m}$ (3.2)
and thus $02\ddot{m}+\kappa\dot{m}-\kappa\dot{m}=$ 0 and this implies that $\ddot{m}=0$ and therefore assuming the boundary conditions that $m(0)=0$ and $m(1)=1$ we have:
$m(t)=t$ (3.3)
The meaning of this is that when a central trading desk will take orders from each trader, assemble them into a single order and then allocate trade fills back to each trader according to the fraction of the total quantity each trader requested. With this we can immediately compute the cost of trading to the central desk.
Proposition 3.1.
If the central trading desk wishes to minimize costs it will trade the risk-neutral strategy $m(t)=t$ and the implementation cost (the aggregate cost) will be the integral of $(\dot{m}+\kappa m)\dot{m}$ and since $\dot{m}=1$ we have
Aggregate $\text{Aggregate cost}=\int_{0}^{1}1+\kappa t\,\mathrm{dt}=1+\frac{\kappa}{2}$ (3.4)
As this represents the implementation cost to the central trading desk, if trader $i$’s target quantity is $\lambda_{i}$ where $\lambda_{i}$ is the fraction of the total quantity being acquired, then trader $i$’s share of the costs will be $\lambda_{i}(1+\frac{\kappa}{2})$. In itself this is interesting but later in Corollary 4.2 be able to give a precise comparison of this cost to the total implementation cost in the case where each trader acts independently and will see that the centralization represents a potentially significant savings. $\blacksquare$
4 The market strategy
In this section we return to the general situation of an $n$-trader non-cooperative game. That is, we assume each of the $n$-traders is merely trying to minimize costs relative to the loss function $C_{i,\lambda_{i}}(t)$. In this case we are able to give a complete description of the equilibrium.
4.1 The market strategy is eager
In this section and the next we show that the market strategy is an eager strategy in the sense of Section 2.3, and use this to calculate the aggregate cost of trading (i.e., the sum of the implementation costs for all traders) has a simple, closed-form solution that is independent of how traders’ target quantities are distributed.
Proposition 4.1.
Let $\lambda_{1},\dots,\lambda_{n}$ be the scaled target quantities of $n$ traders in competition to buy the same stock. If $a_{\lambda_{1}},\dots,a_{\lambda_{n}}$ are a Nash equilibrium then the market strategy $m(t)=\sum_{i=1}^{n}a_{\lambda_{i}}(t)$ is the strategy:
$m(t)=\frac{1-e^{-\alpha t}}{1-e^{-\alpha}},\qquad\alpha=\kappa\frac{n-1}{n+1}$ (4.1)
An immediate corollary of this is that the total implementation cost depends only on $\kappa$ and the number of traders $n$. To see this we only need to note that the total cost is the sum of the individual cost functions. Recalling that aggregate cost is the sum over all $i$ of $C_{i,\lambda_{i}}(t)=\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}\dot{a}_{i}(t)$ and writing
Aggregate $\begin{split}\text{Aggregate cost}&=\sum_{i=1}^{n}\int_{0}^{1}C_{i,\lambda_{i}}(t)\,\mathrm{dt}\\ &=\sum_{i=1}^{n}\int_{0}^{1}\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}\dot{a}_{i}(t)\,\mathrm{dt}=\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\dot{m}(t)\,\mathrm{dt}\end{split}$ (4.2)
A little manipulation of leads to the following immediate corollary whose proof is presented in Section A.3.
Corollary 4.1 (Aggregate cost is independent of target quantities).
Let $a_{\lambda_{i}}$ for $i=1,\dots n$ be an equilibrium as in Proposition 4.1 and let $\text{Cost}(a_{\lambda_{i}})$ be the implementation cost for trader $i$ with respect this equilibrium. Then the aggregate for for all the strategies is given by the expression:
Aggregate $\text{Aggregate cost}=\frac{\kappa\left(\frac{1}{n+1}(e^{-2\alpha}-1)+(1-e^{-\alpha})\right)}{(1-e^{-\alpha})^{2}}=\frac{\alpha}{e^{\alpha}-1}+\frac{\kappa n}{n+1}$ (4.3)
where $\alpha=\kappa\frac{n-1}{n+1}$. Therefore the aggregate cost only depends on the number of traders and the market impact parameters and is linear in the market impact parameter $\kappa$. We note that in Proposition 5.2 we directly derive aggregate cost as the right hand-expression. $\blacksquare$
An immediate consequence of Proposition 4.1 is that for a given value of $\kappa$, the aggregate cost has a limiting value. The total implementation cost has a much simpler form in the limit as the number of traders grows large because $\lim_{n\to\infty}\alpha=\kappa$:
$\lim_{n\to\infty}\text{Aggregate cost}=\frac{\kappa}{e^{\kappa}-1}+\kappa=\frac{\kappa e^{\kappa}}{e^{\kappa}-1}$ (4.4)
Note that the finite $n$ ratio of aggregate cost with our without centralization is given by:
$\left(\frac{\alpha}{e^{\alpha}-1}+\frac{\kappa n}{n+1}\right)\left(1+\frac{\kappa}{2}\right)^{-1}$ (4.5)
This allows us to immediately compute the price of anarchy which is the limiting value as $n\to\infty$ of the ratio of the aggregate cost to the market minimizing cost (see Section 3 $(1+\frac{\kappa}{2})^{-1}$. Figure 3 display the relationship between the price of anarchy, $\kappa$ and the number of traders.
Corollary 4.2.
The price of anarchy in the limit as the number of traders grows large is given by:
$\frac{\kappa}{1-e^{-\kappa}}\left(1+\frac{\kappa}{2}\right)^{-1}<2$ (4.6)
Thus we conclude that the price of anarchy is bounded above by $2$ and this bound is tight. $\blacksquare$
Figure 3: The price of anarchy as it relates to the level alpha-decay and the number of traders. Roughly speaking, the greater the price of anarchy the more the market as a whole benefits from centralization. Equally true is that the greater the $\kappa$, that is the more alpha-decay there is, the more the marfet benefits from centralization.
4.2 Governing differential equations for traders in competition
Sticking with the standard scenario of $n$ traders trading an aggregate quantity of one unit and each trader $i$ trading a portion $10<\lambda_{i}<$ 1 of the aggregate target quantity, we show that if strategies if there a set of scaled strategies $a_{\lambda_{i}}$ (each of which is each equal to $\lambda_{i}a_{i}$ where $a_{i}$ is a unit strategy), $i=1,\dots,n$ are an equilibrium, then they and the resultant market strategy $m=\sum_{i=1}^{n}\lambda_{i}a_{i}$ must satisfy three related differential equations:
Proposition 4.2 (Governing differential equations).
When the trading strategies $a_{\lambda_{i}}$ are an equilibrium, then these strategies along with the market strategy $m$ satisfy the following governing differential equations:
$\ddot{a}_{i}-\kappa\dot{a}_{i}$ $=-\frac{1}{\lambda_{i}}(\ddot{m}+\kappa\dot{m})$ (4.7a)
$\ddot{m}+\kappa\dot{m}$ $=\frac{2\kappa}{n+1}\dot{m}$ (4.7b)
$\ddot{m}+\alpha\dot{m}$ $=0$ (4.7c)
where, as above, $\alpha=\kappa\frac{n-1}{n+1}$ and if it satisfies $m(0)=0,m(1)=1$. An immediate consequence of this is that:
$\ddot{a}_{i}-\kappa\dot{a}_{i}=-\frac{1}{\lambda_{i}}\frac{2\kappa}{n+1}\dot{m}$ (4.8)
The proof Proposition 4.2 is given in Section A. The market strategy $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ (see Eq. 4.1 and therefore substituting this into Eq. 4.7a conclude:
$\ddot{a}_{i}-\kappa\dot{a}_{i}=-\frac{1}{\lambda_{i}}\cdot\frac{2\kappa}{n+1}\cdot\frac{\alpha\,e^{-\alpha t}}{1-e^{-\alpha}}$ (4.9)
This leads to our being able to give a closed form solution for the equilibrium expressions.
5 The equilibrium strategies
In this section compute the equilibrium strategies and explore some of their distinctive properties including the cost of trading in an equilibrium.
Proposition 5.1 (Equilibrium strategies).
Suppose there are $n$ traders labeled $i=1,\dots,n$ trading in competition and that trader $i$’s target quantity is the fraction $10<\lambda_{i}<$ 1 of the total target quantity. Then if each trader $i$ trades the strategy:
$a_{i}(t)=B_{i}(e^{\kappa t}-1)+D_{i}(1-e^{-\alpha t})$ (5.1)
with the constants $B_{i},D_{i}$ and $\alpha$ as follows:
$B_{i}=\frac{\lambda_{i}n-1}{\lambda_{i}n(e^{\kappa}-1)},\qquad D_{i}=\frac{1}{\lambda_{i}n(1-e^{-\alpha})},\qquad\alpha=\kappa\frac{n-1}{n+1}$ (5.2)
and further the resultant market strategy $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ derived by summing the $a_{i}$ in Eq. 5.1 again yields the eager strategy given in Proposition 4.1. In next Section 5.2 we show that for the case that $\lambda_{i}=1/n$ for all $i$ each trader traders a $1/n$-scaled replica of this strategy. Also in proposition 5.2 we will compute the cost of trading each equilibrium strategy.
5.1 Visualizing equilibria
In this section we provide a two of visualizations for the case of three traders to get a feel for what the shape of these strategies looks like. In each plot we show strategies for four different values of $\kappa$, the level of alpha-decay and a fixed distribution of target quantities. All plots scale the trading strategies to a normalized ending quantity of 1 so that we can compare the shapes of each strategy.
In Figure 4 we show three-trader equilibria for the case where traders one to three are respectively 20%, 30% and 50% of the total. We see that as the level of alpha-decay increases from 1 on the left to 20 on the right, strategy 1 with 20% of the total quantity evolves toward a "bucket" strategy, acquiring more than 100% of its target quantity in 20% of the time and then selling (at a profit) toward the end to take advantage of the permanent impact due to the other two traders. We also notice that trader 2 who trades 30% of the total quantity trades a similar strategy but to a lesser extent. Trader three trades a defensive strategy wherein it never overbuys relative to its target quantity.
Figure 4: Each plot shows the same mix of target quantities but for different levels of $\kappa$. All plots show scaled final quantities, for example, while all plots meet at one unit at time 1, in reality strategy one would acquire 20% of the total. Below each plot the implementation cost to each trader is displayed along with the percentage of the total cost it represents. These are calculated using the formula for total cost given in Section A.4. In each panel, the value of $\lambda_{1}$ (trader 1’s percentage of total target quantity) is set to 0.2 and this corresponding strategy is in blue.
Next in Figure 5 we show the same mix of strategies for high alpha-decay.
Figure 5: Each plot shows the same mix of target quantities but for different levels of $\kappa$. All plots show scaled final quantities, for example, while all plots meet at one unit at time 1, in reality strategy one would acquire 20% of the total. Below each plot the implementation cost to each trader is displayed along with the percentage of the total cost it represents. These are calculated using the formula for total cost given in Section A.4. In each panel, the value of $\lambda_{1}$ (trader 1’s percentage of total target quantity) is set to 0.2 and this corresponding strategy is in blue.
5.2 The symmetric equilibrium
An $n$-trader symmetric equilibrium is the case where each trader trades the same quantity of stock and trades the same strategy. This was derived in [Chr24a]. We now show that the general $n$-trader equilibrium contains the symmetric equilibrium as a special case, arrived at from more general principles. Thus we will show that Eq. 4.9 is the same strategy derived in [Chr24a] for the special case $\lambda_{i}=1/n$ and $a_{i}=a_{j}$ for all $i,j$. Plugging these values into the constants for the equilibrium strategies in Eq. 5.2 we obtain $B_{i}=0$ and $D_{i}=\frac{1}{1-e^{-\alpha}}$ so that the equilibrium strategy in Eq. 4.9 becomes:
$a_{i}(t)=\frac{1-e^{-\kappa\frac{n-1}{n+1}t}}{1-e^{-\kappa\frac{n-1}{n+1}}},\qquad\text{for all $i$}$ (5.3)
which precisely recovers the multi-trader symmetric equilibrium in [Chr24a].
5.3 The cost of trading in equilibrium
We can use the facts we have already deduced concerning the market strategy to produce a closed-form expression for the cost of trading to each trader in an equilibrium and gather some information concerning the nature of these costs. We summarize these in the following proposition.
Proposition 5.2 (The cost of trading in equilibrium).
Given $n$ strategies in equilibrium with target quantities $\lambda_{i}$, unit strategies $a_{i}$ and market strategy $m=\sum\lambda_{i}a_{i}$ the cost of trading to trader $i$ $\text{Cost}(\lambda_{i}a_{i},m)$ is as follows:
$\text{Cost}(\lambda_{i}a_{i},m)=\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}$ (5.4)
We give the proof of this in Appendix A.4 but note a few facts here. First, the proof exploits the relationship between $\ddot{a}_{i}-\kappa\dot{a}_{i}$ and $\dot{m}$, without which appears to be very difficult. Second, summing $\text{Cost}(\lambda_{i}a_{i},m)$ over all strategies $i=1,\dots,n$ yields the total implementation cost, or simply aggregate cost. We show in Appendix A.4 that this sum is equal to:
Aggregate $\text{Aggregate cost}=\sum_{i=1}^{n}\text{Cost}(\lambda_{i}a_{i},m)=\frac{\alpha}{e^{\alpha}-1}+\frac{\kappa n}{n+1}$ (5.5)
and while it is not immediately obvious, this is equivalent to the formula for the total implementation cost in Corollary 4.2. We also note that since as $n\to\infty$, $\alpha\to\kappa$, total implementation cost tends to $\frac{\kappa}{e^{\kappa}-1}+\kappa=\frac{1}{1-e^{-\kappa}}$, which is the same as 4.4. $\blacksquare$
We summarize the most important conclusions here:
  • If $n$ traders in competition are in equilibrium, the there are only three variables that contribute to implementation cost: (a) the fraction of the total quantity traded by the trader (e.g., for trader $i$ this is $\lambda_{i}$), (b) the total number of traders, and (c) the market impact parameter.
  • We can see from Eq. 5.4 that as a given trader’s share, say $\lambda_{i}$, approaches 100% of the cost, but the limit is complicated by the fact that in this case the number of traders approaches $1$.
5.4 Each trader’s implementation as a share of the total
We have already seen that the total implementation costs over all traders does not depend on the distribution of target quantities, despite the fairly large differences between the shapes of each strategy. Therefore it’s reasonable to ask whether each traders share of total cost is simply the their expected share of the total cost. For example, if a trader is trading 25% of the total quantity it might be the case that their share of the total cost is simply 25% of the total. We answer this firmly in the negative in Proposition 5.3.
Each trader’s share of the total implementation cost is of interest because we’d like to know under what circumstances a trader is paying more or less than their expected share of total costs. The implementation cost to trader $i$ is Eq. 5.4 and the total implementation cost is Eq. 5.5 and thus the ratio is:
$S(\lambda_{i}a_{i};\kappa,n)=\frac{\text{Cost}(\lambda_{i}a_{i},m)}{\text{Aggregate cost}}.$ (5.6)
We summarize the ratio’s formula in the following Proposition.
Proposition 5.3.
Given the equilibrium strategies $\{\lambda_{1}a_{1},\dots,\lambda_{n}a_{n}\}$ with alpha-decay parameter $\kappa$, the share of trader $i$’s implementation cost is given by:
$S(\lambda_{i}a_{i};\kappa,n)=\frac{1}{n}+\frac{(n+1)(e^{\alpha}-1)}{n(1-e^{-\kappa})(1-ne^{\alpha})}-\frac{(n+1)(e^{\alpha}-1)}{(1-e^{-\kappa})(1-ne^{\alpha})}\lambda_{i}$ (5.7)
where as usual $\alpha=\kappa\frac{n-1}{n+1}$. This is clearly linear in $\lambda_{i}$ meaning that any specific trader’s implementation cost as a percentage of total cost increases in proportion to their share. The proof is given in Appendix A.5 and we provide some visualizations next. $\blacksquare$
5.5 Visualizations of implementation cost share
The implementation cost share is a measure of the fraction of the total cost that a trader pays relative to what one might think of as the expected cost, namely since trader $i$’s share of the target quantity is $\lambda_{i}$ then one might expect that the traders share of the aggregate cost might be $\lambda_{i}$ as well. However this is note the case.
Figure 6 provides a visualization of this phenomena. The plots show a comparison between the actual costs and fair-share costs. Each plot is set up as follows:
  • For a trader (say 1) with a target quantity equal to $\lambda_{1}$, where this represents the fraction of the total quantity, the fair-share cost is $\lambda$. For example, if for a given trader $\lambda_{1}=0.2$ (20% of the total quantity, then their fair-share cost is 20%.
  • Each plot represents a specific number of trader $n$ (e.g., the upper-left has $n=2$, two total traders) and a given $\kappa$ (e.g., the upper-left has $\kappa=0.5$).
  • The bars in each plot show the deviation between the fair-share cost and actual percentage of total cost in equilibrium for trader 1 for different percentage of total target quantity, denoted $\lambda_{1}$. For example, in the upper-left plot $\lambda_{1}=0.01$ means for the case where trader 1’s target quantity is 1% of the total, when there are total two traders and $\kappa=0.5$, trader 1 pays less than its fair share by approximately 0.6%. Meaning, roughly, that instead of 1% of the total cost, they pay approximately 0.4%.
Figure 6: The deviation of a given trader’s share of implementation cost from "fair share", where fair share means the same fraction of total cost as fraction of trade. For example, if trader 1 is 20% of the trade (i.e., $\lambda_{1}=0.2$), then their fare share is 20%. The bars in each plot show the change (up or down) from fair share. For example, the lower right plot shows that for traders trading 88% and 99% of the total target quantity, the share they actually pay is more than 4% greater than fair share. This means that the trader trading 99% will pay more than 100% of total cost (and thus some traders earn a profit from the trade.
Note that in Figure 6 the share of total cost for a given trader can exceed 100%. For example, in the upper-right plot where $n=2$ (two traders) and $\kappa=10$ the bar for trader 1 with $\lambda_{1}=0.99$ is a little more than +20%. This means that trader 1 in this case pays over 120% of the total cost.
6 Centralized trading
We now study the question can a firm that employs a portion of all traders in competition to buy a stock lower its implementation costs by centralizing its traders? The short answer is, it depends, that is, the details of what precisely the firm does with the centralized trades. We characterize that details of what we mean by trade centralization:
To fix standard terminology, our setup will be that there are $n$ total traders and $n_{1}$ of them work for a firm, which we will referred to as simple "the firm" with traders "firm traders", while the remaining $n_{2}$ traders work independently and referred to as "the non-firm". As usual traders are labeled $i=1,\dots,n$ and each trader’s target quantity’s portion of the total is $\lambda_{i}$. We write $\lambda_{n_{1}}=\sum_{i=1}^{n_{1}}\lambda_{i}$ for the fraction of the total quantity traded by the firm, and $\lambda_{n_{2}}=\sum_{i=n_{1}+1}^{n_{2}}\lambda_{i}$ for the fraction traded by the non-firm traders:
$\underbrace{\lambda_{1},\dots,\lambda_{n_{1}}}_{\text{Firm, $\lambda_{n_{1}}$}}\quad\text{and}\quad\underbrace{\lambda_{n_{1}+1},\dots,\lambda_{n_{2}}}_{\text{Non-firm, $\lambda_{n_{2}}$}}$ (6.1)
Each trader within the firm has a target quantity and associated parent order associated they would like to execute (see Section 2.2. If the firm funnels all the parent orders through an internally operated trading desk, the central trading desk, then it is said to be centralizing its trades. A prominent reason for trade centralization is to reduce implementation costs. The thinking behind this is that if each trader trades independently then they will operate to minimize their own implementation costs.
As we saw in Section 3 the market as a whole reduces its costs relative to all traders fending for themselves. Therefore it seems likely that centralizing trades would benefit the firm through lower aggregate implementation costs. In truth, it is not so simple. It turns out that whether or not a firm can lower its costs through centralization hinges on exactly what the firm does with the aggregated parent orders. As this is of considerable interest, we first carefully define some terms.
  • Firm order flow: this is the firms detailed plan for implementing the aggregate of all firm trader’s parent orders with respect to the stock they are competing to purchase;
  • Non-cooperative trading: this refers to the situation where all traders are trading without coordinating or communicating with others. This term is with reference to non-cooperative games in which players are not coordinating or communicating;
  • Cooperative trading: when a set of traders implicitly or explicitly pool their orders and cooperate to reduce their aggregate implementation costs. This is implicitly the case when a firm centralizes its trades and seeks to minimize firm costs which is tantamount to minimizing aggregate costs;
Given this setup there are two possible approaches the firm can take. The first case is that they can allow each trader to trade independently (the non-cooperative approach) and the second is the firm can centralize all the trades (the cooperative approach). We summarize as in the following table:
Firm Non-firm
Strategy Traders Traders Quantity Traders Quantity
Non-cooperative $n$ $n_{1}$ $\lambda_{n_{1}}$ $n_{2}$ $\lambda_{n_{2}}$
Cooperative $n_{2}+1$ 1 $\lambda_{n_{1}}$ $n_{2}$ $\lambda_{n_{2}}$
Table 1: Firm and non-firm trading setup as it pertains to how the parameters of competitive trading change when moving from cooperative to non-cooperative trading.
In the following sections we aggregate the strategies in each row of Table 1 into two consolidated strategies: the firm strategy and the aggregate non-firm strategy. Like the market strategy these are not always strategies, however we treat them as such so that we can readily refer to their costs and study their properties.
With these definitions out of the way, we distinguish between the two possible modes of centralization we will discuss. In both cases the basic setup is as in
  • Naive centralization: when a firm naively centralizes it merely takes all of the order flow and treats it as a single trade. Then the game transitions from an $n$-trader game to an $n_{2}+1$ trader game where $n_{2}$ is the number of non-firm traders and 1 represents the aggregated order flow of the firm. We call this naive because the firm trades the aggregated parent orders as one without regard to other strategic options; and
  • Strategic centralization: in strategic centralization the firm seeks strategic options for what to do with the consolidated order flow. It does not necessarily trade it as a single strategy but seeks the optimal re-organization of the flow that minimizes implementation cost.
6.1 Naive centralization
When a firm naively centralizes it implicitly takes the view that the best way to minimize implementation costs is to trade the centralized order flow as a single monolithic strategy designed to minimize its total implementation cost. In this section, then, we study what this strategy looks like in equilibrium and compare its implementation cost to that of the de-centralized strategies (also in equilibrium). The parameters of the trading game immediately become those of the cooperative row of Table 1. Given this, there are two instances of the firm strategy: one when the traders trade independently and the other where their trades are centralized. We refer to these as the non-cooperative and cooperative strategies respectively.
[…]
$\lambda_{n_{1}}a_{n_{1}}$ $=\lambda_{n_{1}}(B_{n_{1}}(e^{\kappa t}-1)+D_{n_{1}}(1-e^{-\hat{\alpha}t})),\qquad\text{\em Cooperative firm strategy}$ (6.2b)
with the constants $B_{i}$ and $D_{i}$ as in Eq. 5.2:
$B_{i}$ $=\frac{\lambda_{i}n-1}{\lambda_{i}n(e^{\kappa}-1)},\quad D_{i}=\frac{1}{\lambda_{i}n(1-e^{-\alpha})}$ (6.3a)
$B_{n_{1}}$ $=\frac{\lambda_{n_{1}}(n_{2}+1)-1}{\lambda_{n_{1}}(n_{2}+1)(e^{\kappa}-1)},\quad D_{n_{1}}=\frac{1}{\lambda_{n_{1}}(n_{2}+1)(1-e^{-\hat{\alpha}})}$ (6.3b)
and $\alpha=\kappa\frac{n-1}{n+2}$, $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+1}$ and $n_{1}+n_{2}=n$. Next there are the corresponding strategies for the non-firm traders:
[…]
where $B_{i}$ and $D_{i}$ are as above and $\hat{B_{i}}$ and $\hat{D_{i}}$ are as follows:
$\hat{B}_{i}=\frac{\lambda_{i}(n_{2}+1)-1}{\lambda_{i}(n_{2}+1)(e^{\kappa}-1)},\quad\hat{D}_{i}=\frac{1}{\lambda_{i}(n_{2}+1)(1-e^{-\alpha})}\$ (6.5)
In Figure 7 we display comparisons of the firm and non-firm strategies in the cooperative versus the non-cooperative cases. In each plot green lines represent equilibrium aggregated strategies when firm traders centralize their trades (the cooperative case) and blue lines represent the same but for the non-cooperative (when traders do not centralize). Solid lines represent the firm’s aggregated strategy in equilibrium and dashed lines represent the non-firm traders’ aggregated strategy. We can think of a pair of lines of the same color (one solid and one dashed) as a strategy and a counter-strategy. We now discuss the meaning of this.
Figure 7: A comparison of aggregated firm and non-firm strategies with and without trade centralization. In this case there are six total traders, and from left to right the number of traders in the firm is 2, 3 and 5 respectively and the firm’s quantity as a fraction of the total grows from 50% to 95%.
First, since dashed lines represent the non-firm traders’ aggregate strategy, comparing dashed green to dashed blue lines is comparing the shape of non-firm strategies when the firm centralizes (the cooperative case, in green) versus when it doesn’t. It is clear that in this case cooperation results in the non-firm traders trading more aggressively and therefore front-running to a greater degree (see Section 2.3). But what happens to the firm’s strategy when trades are centralized? In this case, the firm represents a greater share of the total target quantity and the total number of traders, the firm’s strategy becomes generally more defensive and the result is it is more readily front-run (see Section 2.3).
Figure 8 we provide a second example of the how centralization impacts trades.
Figure 8: Another example of firm versus non-firm aggregated strategies in cooperative (centralized) versus non-cooperative situations, see Figure 7.
Figures 7 and 8 suggest a surprising hypothesis. Firms that consolidate trades may not be gaining by doing so. We examine this further next.
6.2 Naive centralization implementation costs
Continuing with the same setup as above, we will explicitly compute the cost of each of the four strategies above, that is for the aggregated firm and non-firm strategies for the cooperative and non-cooperative scenarios in Figures 7 and8.
We will compare the costs of each approach. First compute the non-cooperative firm implementation cost, specifically this covers the case when traders $1,\dots,n_{1}$ trade independently of one another. This is nothing more than the sum of the costs in (5.4):
Firm cost ($\begin{split}\textit{Firm cost (non-cooperative)}&=\sum_{i=1}^{n_{1}}\text{Cost}(\lambda_{i}a_{i},m)\\ &=\sum_{n=1}^{n_{1}}\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}\\ &=\kappa\frac{\lambda_{n_{1}}n-n_{1}}{n(1-e^{-\kappa})}+\alpha\frac{n_{1}}{n(e^{\alpha}-1)}+\frac{n_{1}\kappa}{n+1}\end{split}$ (6.6)
where as usual $\alpha=\kappa\frac{n-1}{n+1}$. We can similarly compute the non-firm traders’ cost in the non-cooperative case:
Non-firm cost ($\begin{split}\textit{Non-firm cost (non-cooperative)}&=\sum_{i=n_{1}+1}^{n}\text{Cost}(\lambda_{i}a_{i},m)\\ &=\sum_{n=n_{1}+1}^{n}\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}\\ &=\kappa\frac{\lambda_{n_{2}}n-n_{2}}{n(1-e^{-\kappa})}+\alpha\frac{n_{2}}{n(e^{\alpha}-1)}+\frac{n_{2}\kappa}{n+1}\end{split}$ (6.7)
where $\lambda_{n_{2}}=\sum_{i=n_{1}+1}^{n}\lambda_{i}$ represents the fraction of the total target quantity sought be the non-firm traders. Now we can compute the same quantities in the cooperative case. On the other hand, when the firm trades are centralized, the firm trades the cooperative firm strategy in which case the total number of traders is $n_{2}+1$ consisting of the $n_{2}$ non-firm traders and the firm. The firm’s strategy is labeled $a_{n_{1}}$. The cost of implementing the strategy $\lambda_{n_{1}}a_{n_{1}}$ is calculated applying Eq. 5.4 setting the number of traders to $n_{2}+2$ and replacing $\alpha$ with $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+2}$:
[…]
Next we calculate the non-firm traders’ aggregate implementation cost when firm centralizes its trades. In this case the total number of traders is $n_{2}$ and we set $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+2}$ as above and sum over the implementation cost of each strategy:
Non-firm costs ($\begin{split}\textit{Non-firm costs (cooperative)}&=\sum_{i=n_{1}+1}^{n}\kappa\frac{\lambda_{i}(n_{2}+1)-1}{(n_{2}+1)(1-e^{-\kappa})}+\hat{\alpha}\frac{1}{(n_{2}+1)(e^{\hat{\alpha}}-1)}+\frac{\kappa}{n_{2}+2}\\ &=\kappa\frac{\lambda_{n_{2}}(n_{2}+1)-n_{2}}{(n_{2}+1)(1-e^{-\kappa})}+\hat{\alpha}\frac{n_{2}}{(n_{2}+1)(e^{\hat{\alpha}}-1)}+\frac{\kappa n_{2}}{n_{2}+2}\end{split}$ (6.9)
6.3 Detailed numerical examples of naive centralization
We now provide a detailed examination of the difference in cost when a firm centralizes its trades based on the implementation cost formulas presented above. As in the prior section there is a firm with $n_{1}$ traders, $n_{2}$ other traders and (the non-firm traders) and the total target quantity is scaled to one. The firm’s portion of the target quantity is $\lambda_{n_{1}}$. This section provides data and analysis of the change in implementation costs to the firm and the non-firm traders in the aggregate, when the firm switches to centralized trading. Note that the general picture may seem counter-intuitive, as firm’s do not always or generally experience lower implementation costs (at least as far as the models go) when they centralize, because as we will see the firm becomes a target for front-running as its trades become larger and more predictable (see Section 2.3). Note also that in Section 6.4 we look at at related and more general scenarios where a firm disguises the number of traders it actually has in order to reduce implementation costs.
Tables 2 and 3 display the average change in implementation cost when a firm switches all of its traders from trading independently to trading through a centralized desk. The reason we say average here is because each row in each table is an average over randomly selected values for the distribution of target quantities, number of firm traders and total number of traders. The two tables are split according to the average number of traders in the firm relative to non-firm traders. In both tables the total number of traders ($n$) is randomly selected from 20, 21 or 22, but in the first the number of firm traders ($n_{1}$) is randomly selected from 3, 4 or 5, while in the second it is selected from 14, 15 or 16. Each table is broken up into three blocks according to the value of $\kappa$, representing the level of alpha-decay. Each row then is an average value of the
The big picture may be summarized as when a firm centralizes its trades the total implementation over all traders decreases but the benefits accrue to the non-firm traders and not the firm. As this is undoubtedly a surprising result it’s important to pin down precisely what the underlying assumptions are. For starters, for this to hold the market must realize that the firm is trading as a single monolithic block and adjust to this to form a new equilibrium. With this said, we make the following comments regarding this. When a firm centralizes its trades the following occur:
  • The aggregate firm implementation cost generally increases especially when the number of traders in the firm is a high proportion of all traders;
  • Overall (across all traders firm and non-firm) implementation costs decrease, consistent with the market minimization results in Section 3. However, these improvements generally accrue to the non-firm traders;
  • The implementation cost increase to the firm is greatest on a percentage basis when the firm is trading a small fraction of the total quantity.
Percent Change Costs (Firm) Costs (Non-firm)
$\lambda_{n_{1}}$ $n_{1}$ $\kappa$ Firm Non-firm Total Non-Coop Coop Non-Coop Coop
0.07 4.00 1.00 1.6% -0.3% -0.1% 0.12 0.12 1.45 1.45
0.15 4.00 1.00 0.8% -0.3% -0.1% 0.23 0.24 1.33 1.33
0.40 4.00 1.00 0.3% -0.4% -0.1% 0.63 0.63 0.94 0.93
0.62 4.00 1.00 0.2% -0.7% -0.1% 0.99 0.99 0.58 0.58
0.82 4.00 1.00 0.1% -1.7% -0.1% 1.30 1.30 0.27 0.26
0.07 4.00 5.00 8.7% -1.3% -0.7% 0.34 0.36 4.49 4.43
0.15 4.00 5.00 4.0% -1.5% -0.7% 0.71 0.74 4.11 4.05
0.40 4.00 5.00 1.4% -2.2% -0.7% 1.97 2.00 2.85 2.79
0.62 4.00 5.00 0.9% -3.8% -0.7% 3.11 3.13 1.72 1.66
0.82 4.00 5.00 0.7% -10.1% -0.7% 4.11 4.14 0.71 0.65
0.07 4.00 25.00 9.5% -1.5% -0.8% 1.66 1.80 22.20 21.88
0.15 4.00 25.00 4.4% -1.6% -0.8% 3.53 3.68 20.33 20.00
0.40 4.00 25.00 1.5% -2.4% -0.8% 9.78 9.93 14.08 13.75
0.62 4.00 25.00 1.0% -4.2% -0.8% 15.41 15.55 8.45 8.13
0.82 4.00 25.00 0.7% -11.3% -0.8% 20.41 20.55 3.45 3.13
Table 2: The average value of change in implementation costs for a variety of scenarios when a firm switches from non-cooperative trading to centralized trading. In each case the total number of traders is one of 20, 21 and 22 and the average number of traders in the firm in this table is always 4, meaning the traders represent a minority of all traders. Each block of rows is a variety of scenarios for a fixed $\kappa$ and what varies is the fraction of the total target quantity being that the firm represents, denoted as $\lambda_{1}$. The percent changes columns display the change in implementation cost switching for the firm, for the aggregate of the non-firm and firm traders.
Percent Change Costs (Firm) Costs (Non-firm)
$\lambda_{n_{1}}$ $n_{1}$ $\kappa$ Firm Non-firm Total Non-Coop Coop Non-Coop Coop
0.07 15.00 1.00 4.6% -1.8% -1.4% 0.11 0.11 1.46 1.43
0.15 15.00 1.00 2.2% -2.0% -1.4% 0.23 0.23 1.34 1.31
0.40 15.00 1.00 0.8% -2.9% -1.4% 0.62 0.63 0.95 0.92
0.62 15.00 1.00 0.5% -4.8% -1.4% 0.98 0.98 0.59 0.56
0.82 15.00 1.00 0.4% -10.9% -1.4% 1.30 1.30 0.27 0.25
0.07 15.00 5.00 6.8% -1.5% -1.4% 0.30 0.31 4.55 4.51
0.15 15.00 5.00 3.1% -1.7% -1.4% 0.64 0.66 4.19 4.15
0.40 15.00 5.00 1.0% -2.6% -1.4% 1.79 1.81 2.87 2.83
0.62 15.00 5.00 0.6% -4.4% -1.4% 2.95 2.97 1.73 1.69
0.82 15.00 5.00 0.4% -10.7% -1.4% 3.94 3.96 0.74 0.70
0.07 15.00 25.00 9.0% -1.4% -1.5% 1.57 1.72 22.22 21.91
0.15 15.00 25.00 4.2% -1.6% -1.5% 3.37 3.53 20.36 20.05
0.40 15.00 25.00 1.4% -2.3% -1.5% 9.61 9.77 14.09 13.78
0.62 15.00 25.00 0.9% -4.1% -1.5% 15.23 15.39 8.46 8.16
0.82 15.00 25.00 0.6% -11.2% -1.5% 20.22 20.38 3.46 3.16
Table 3: The average value of change in implementation costs for a variety of scenarios when a firm switches from non-cooperative trading to centralized trading. In each case the total number of traders is one of 20, 21 and 22 and the average number of traders in the firm in this table is always 15, meaning the traders represent a majority of all traders. Each block of rows is a variety of scenarios for a fixed $\kappa$ and what varies is the fraction of the total target quantity being that the firm represents, denoted as $\lambda_{1}$.The percent changes columns display the change in implementation cost switching for the firm, for the aggregate of Independents and in the aggregate.
6.4 Strategic centralization
Given the observations above, namely that centralizing trades counter-intuitively increases implementation costs, it stands to reason that reversing the process, namely of de-centralizing trades in some manner might decrease costs. Can a firm even go further and disguise the true number of traders it has to fool the market into believing it has more than it does. We call this process disguising order flow and examine in what circumstances and to what extent it can improve implementation costs.
examine the change in implementation costs when a firm parcels out its trades in order to appear as if the number of non-firm traders in the market is different than it actually is. We assume we are in the same situation as when studying trade centralization (see Section 6). There is a firm with $n_{1}$ traders whose target quantity in the aggregate is $\lambda_{n_{1}}$ and there are $n_{2}$ traders outside of the firm (the non-firm traders) whose target quantities in the aggregate are $\lambda_{n_{2}}$. The total number of traders is $n=n_{1}+n_{2}$.
[…]
Now imagine that the wants to "pretend" that the firm sends its orders out in such a way as change the number of traders from $n_{1}$ to $n_{1}+m$ where $m+n_{1}\geq 1$ so that the firm masks the real number of traders it has and makes it appear as if there are $n_{1}+m$ so that if $m<0$ it is reducing the number of trades and if $m>0$ it is increasing the number of trades. In all cases the firm still has a total target quantity of $\lambda_{n_{1}}$. Thus the firm has disguised its true number of traders by altering the umber of traders. In this case we have:
$n_{1}\to n_{1}+m,\quad n\to n+m\quad\alpha\to\hat{\alpha}=\kappa\frac{n+m-1}{n+m+1}$ (6.11)
and therefore firm aggregate implementation cost when the number of traders goes from $n_{1}$ to $n_{1}+m$:
[…]
Note that in theory we can take the derivative of Eq. 6.12 with respect $m$ and calculate where its zeros are in order to see at what value of $m$ costs are minimized, however we will only explore this numerically. One thing that we can do, however, is let $m\to\infty$ we see that $\hat{\alpha}\to\kappa$ and $(n_{1}+m)/(n+m)\to 1$ therefore the first term in Eq. 6.10 tends to $\frac{\kappa\lambda_{n_{1}}-\kappa}{1-e^{-\kappa}}$, while the second and third terms tend to $\frac{\kappa}{e^{\kappa}-1}$ and $\kappa$ respectively. Therefore as $m\to\infty$ the firm implementation cost tends to a limiting value of:
$\begin{split}\frac{\kappa\lambda_{n_{1}}-\kappa}{1-e^{-\kappa}}+\frac{\kappa}{e^{\kappa}-1}+\kappa&=\frac{\kappa\lambda_{n_{1}}-\kappa}{1-e^{-\kappa}}+\frac{\kappa e^{-\kappa}}{1-e^{-\kappa}}+\kappa\frac{1-e^{-\kappa}}{1-e^{-\kappa}}\\ &=\frac{\kappa\lambda_{n_{1}}}{1-e^{-\kappa}}\end{split}$ (6.13)
We can perform a similar procedure for the non-firm trader’s implementation costs where $n_{2}$ is the number of traders not in the firm. Here the situation is particularly simple because the $n_{2}/n\to 0$ as the number of splits grows large. Therefore the non-firm traders’ costs tends to
[…]
and therefore in the limit as the number of traders grows large we have the non-firm traders’ (non-firm traders) aggregate implementation costs tends to $\frac{\kappa\lambda_{n_{2}}}{1-e^{-\kappa}}$ as the number of traders grows large,
6.5 Visualizations of strategic centralization implementation costs
We provide visual analysis of the impact of implementation costs when a firm disguises the true number of traders to make it appear as if there are more traders than there actually are. We adopt the conventions of the previous section and letting $n_{1}$ be the actual number of firm traders, $n$ be the total number of traders. Below are a number of plots grouped into a few categories. The format is similar throughout. Each plot plots the number $m$ on the $x$-axis and the percentage change in implementation costs versus the baseline of the actual number of traders. This change is naturally the change that would occur if all traders traded a new equilibrium based on the apparent number of traders. In other words, we assume the subterfuge works.
All plots have three lines representing three values for the fraction the firms total quantity is relative to the total quantity, represented by $\lambda_{n_{1}}$=10(the blue lines), $\lambda_{2}$=33.3(the orange lines) and $\lambda_{3}=$ 66.6% (the green lines) of the total quantity. In cases where implementation costs rise upon adding any number of traders, we also plot the reverse where the number of traders is disguised to appear as if there are fewer traders than there actually are.
The firm has one trader and is a large fraction of all traders: we make three plots (each with three sub-plots) in which the firm always has one actual trader ($n_{1}=1$) and the total number of traders is 2, 3 and 10 so that the firm is 1/2, 1/3 and 1/10th of the total number of traders. We then examine implementation cost changes for three levels of alpha-decay, $\kappa=1,2.5$ and $5$. There are a few obviously interesting trends.
Firm is most of the traders: Below we see the case where the firm can substantially benefit from reverse disguising and making it appear as if there are few traders than the actually are. In these plots the black dot represents the baseline in which the firm has not tried to disguise the number of traders.
The firm has one trader in a crowded trade: In the next set of plots the firm has only one trader. The question is whether or not the firm can benefit from disguising its order flow to lower costs. The answer has similar shape to the examples above. As in many, all of the gains occur when adding only a few traders and then very rapidly the gains dissipate and ultimately costs increase.
The firm represents two-thirds of all traders: in the plots below the firm’s total number of actual traders varies: $n_{1}=10,20,50$ from left to right respectively and in all cases the total number of traders is 3/2 of the firm’s so that the firm is 2/3rds of the total. In these cases adding traders generally raises implementation costs and the firm benefits from making it appear as if it has fewer traders than it actually does. In the plots below, negative values in the x-axis represent when the firm has consolidated trades and the black dot is the baseline when the firm has not disguised its number of traders
6.6 What is really going on with strategic centralization?
The conclusion of the prior sections on centralization is that a firm can improve its implementation costs, but it requires an artful masking of the true number of traders to optimally minimize costs. The question is what is really going on and how is this possible at all?
The answer relates to the fact that when a firm changes the apparent number of traders in the market, they are changing the game that all traders in competition are playing. Recall that the aggregate cost of trading in Eq. 5.5 is an increasing function of the number of traders $n$, while a trader’s cost share given by 5.7 has a complicated relationship in the number of traders $n$. We explore this now to gain intuition regarding internalization.
To start, note that as a system the total cost of trading varies depending on the number of traders. So when a firm masks its true number of traders it is in effect changing the game that is being played itself. Ordinarily this is not possible because each trader is totally independent of all the others. However, when there is a firm involved and that firm employs a number of traders then the firm can effectively modify the number of traders in the market to its advantage. Exactly what that advantage is depends on two opposing forces.
The first of those forces is that as the number of traders in competition increases, all else equal the total cost of trading increases as well. This is a consequence of the degradation of efficiency a system undergoes in a non-cooperative game. The second is that for a given trader, their share of the total cost changes when the number of traders changes. When the number of traders is very small (two or three, say) then increasing the number of traders increases share of cost for a given trader and this increase is significantly greater for traders seeking a large fraction of the total quantity. On the other hand, for a moderate number of traders, the reverse of true and increasing traders decreases cost share.
Figure 9: Each plot shows the percentage change in total implementation cost (over all traders) as the number of traders increases from 2 to 25 where the change is relative to the two-trader case. The plots show that as more traders compete, the total cost to all traders grows and that the extent of this growth is greater as alpha-decay (given by $\kappa$) grows. For example, when $\kappa=25$, the total cost is over 40% higher for 25 traders than 2 traders in competition. On the other hand, for a given trader, the right-hand $y$-axis shows the change in their share of the total cost that occurs when adding one new trader. A given point on the plot represents the change in cost share to a given trader when one more trader is added. One can see that when the number of traders is small, adding a trader increases costs and removing a trader therefore decreases costs.
To unpack Figure 9, note a few things about its content.
  • Each plot is for a different level of alpha-decay given by $\kappa$;
  • Each plot has plots two types of quantities as they relate to the total number of traders $n$, shown on the $x$-axis.
  • The solid lines are the change in the total cost of trading relative to when there are two traders, as a function of the total number of traders using the formula Eq. 5.5. For example, at $n=10$, the plot displays a little more than 3% for the case $\kappa=1.0$, meaning that relative to 2 traders the total cost of trading for 10 traders is 3% higher;
  • The dashed lines are plotted on the right-hand axis and represent the change in a trader’s share of total cost that occurs in adding one more trader to the market. There are four lines in each plot representing four different traders with share of total quantity ranging from 20% to 90%. We can see in the right-hand plot ($\kappa=25$) that traders 3 and 4 decrease costs by 2% when one more trader is added to the market.
  • Note that when the number of traders is relatively small, total cost is most rapidly increasing in number of traders, and at the same time for a moderate but not very small number of traders, increasing the number of traders decreases costs. This is the specific reason that there is a "sweet spot" for trade centralization in terms of number of traders.
The totality of these results points to an explanation of the counter-intuitive features of trade centralization.
7 Acknowledgements
I extend my gratitude to the Machine Learning Research Group at Morgan Stanley and Vlad Ragulin who provided much discussion and feedback. I would also like to thank Michael Kearns and Mirah Shih of the University of Pennsylvania for extremely valuable discussions concerning algorithmic game theory.
Appendix A Proofs of Proposition 4.1 and 5.1
We present a of Proposition 4.1, which says that the market strategy is an eager strategy with a well-defined form.
A.1 Proof of Proposition 4.1
To prove the equality in Eq. 4.7a we use the Euler-Lagrange equation on the loss function $C_{i,\lambda_{i}}=(\dot{m}+\kappa m)\lambda_{i}a_{i}$. First we expand the loss function:
$C_{i,\lambda_{i}}=\left(\sum_{j}\lambda_{j}\dot{a}_{j}+\kappa\sum\lambda_{j}a_{j}\right)\lambda_{i}\dot{a}_{i}$ (A.1)
and then compute the partial derivatives:
$\begin{split}\frac{\partial}{\partial t}\frac{\partial C_{i,\lambda_{i}}}{\partial\dot{a}_{i}}&=\sum_{j\neq i}^{n}\lambda_{i}\lambda_{j}\ddot{a}_{j}+2\lambda_{i}^{2}\ddot{a}_{i}+\kappa\sum_{j=1}^{n}\lambda_{i}\lambda_{j}\dot{a}_{j}\\[5.0pt] \frac{\partial C_{i,\lambda_{i}}}{\partial a_{i}}&=\kappa\lambda^{2}_{i}\dot{a}_{i}\end{split}$ (A.2)
The Euler-Lagrange equations says that $\frac{\partial}{\partial t}\frac{\partial C_{i,\lambda_{i}}}{\partial\dot{a}_{i}}-\frac{\partial C_{i,\lambda_{i}}}{\partial a_{i}}=0$ which implies:
$\sum_{j=1}^{n}\lambda_{i}\lambda_{j}\ddot{a}_{j}+\lambda_{i}^{2}\dot{a}_{i}+\kappa\sum_{j=1}^{n}\lambda_{i}\lambda_{j}\dot{a}_{j}-\kappa\lambda_{i}^{2}\dot{a}_{i}=0$ (A.3)
Dividing everything by $\lambda_{i}$ substituting $\ddot{m}=\sum_{i=1}^{n}\lambda_{i}\ddot{a}_{\lambda_{i}}$ and $\dot{m}=\sum_{i=1}^{n}\lambda_{i}\dot{a}_{\lambda_{i}}$ we obtain:
$\ddot{m}+\lambda_{i}\ddot{a}_{i}+\kappa\dot{m}-\kappa\lambda_{i}\dot{a}_{i}=0$ (A.4)
and rearranging we arrive at:
$\lambda_{i}(\ddot{a}_{i}-\kappa\dot{a}_{i})+\ddot{m}+\kappa\dot{m}=0\implies\ddot{a}_{i}-\kappa\dot{a}_{i}=-\frac{1}{\lambda_{i}}(\ddot{m}+\kappa\dot{m})$ (A.5)
and this proves the equality Eq. 4.7a. To complete the proof we take the weighted-sum of the right-hand-side of Eq. A.5 with weights equal to $\lambda_{i}$ to conclude:
$\ddot{m}-\kappa\dot{m}=-n(\ddot{m}+\kappa\dot{m})$ (A.6)
rearranging this prove the equality Eq. 4.7c $\blacksquare$
A.2 Proof of Proposition 5.1
We seek to show that the functions $a_{i}$ stated in Eq. 5.1 satisfy Eq. 4.7a, which states:
$\ddot{a}_{i}-\kappa\dot{a}_{i}=-\frac{1}{\lambda_{i}}(\ddot{m}+\kappa m)$ (A.7)
To prove this we start by computing the derivatives of $a_{i}$:
$\begin{split}\dot{a}_{i}&=\kappa B_{i}e^{\kappa t}+\frac{\alpha e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}\\ \ddot{a}_{i}&=\kappa^{2}B_{i}e^{\kappa t}-\frac{\alpha^{2}e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}\\ \end{split}$ (A.8)
Then forming $\ddot{a}_{\lambda_{i}}-\kappa\dot{a}_{\lambda_{i}}$ the terms with $B_{i}e^{\kappa t}$ cancel out and we end up with
$\begin{split}\ddot{a}_{\lambda_{i}}-\kappa\dot{a}_{\lambda_{i}}&=-\frac{\alpha^{2}e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}-\kappa\frac{\alpha e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}\\ &=-\left(\frac{\alpha^{2}e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}+\kappa\frac{\alpha e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}\right)\\ &=-\frac{\alpha e^{-\alpha t}}{\lambda_{i}n(1-e^{-\alpha})}(\alpha+\kappa)\end{split}$ (A.9)
Now note that
$\alpha+\kappa=\kappa\frac{n-1}{n+1}+\kappa=\kappa\frac{2n}{n+1}$ (A.10)
plugging this into the last line above we obtain
$\ddot{a}_{\lambda_{i}}-\kappa\lambda_{i}\dot{a}_{i}=-\frac{2\cdot\kappa\cdot\alpha e^{-\alpha t}}{\lambda_{i}(n+1)(1-e^{-\alpha})}$ (A.11)
as desired. $\blacksquare$
A.3 Proof of the total implementation cost equations
We now prove Corollary 4.1 which gives a closed-form formula for the aggregate cost, that is, the aggregate implementation cost. Given $n$ strategies in competition
$\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\dot{m}(t)\,\mathrm{dt}$ (A.12)
where $m(t)$ is the market strategy, $\sum_{i=1}^{n}a_{\lambda_{i}}$. Proposition 4.1 shows:
$m(t)=\frac{1-e^{-\alpha t}}{1-e^{-\alpha}}$ (A.13)
and the cost of trading for each trader $i$ is $\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\dot{a}_{\lambda_{i}}\,\mathrm{dt}$, and summing across all traders shows that the aggregate of the implementation costs over all traders:
Aggregate $\text{Aggregate cost}=\int_{0}^{1}(m(t)+\kappa\,\dot{m}(t))\dot{m}(t)\,\mathrm{dt}$ (A.14)
So to obtain a closed form solution we first need to compute the integrand explicitly. To start note that:
$\dot{m}+\kappa m=\frac{-\frac{2\kappa}{n+1}e^{-\alpha t}+\kappa}{1-e^{-\alpha}}$ (A.15)
and $\dot{m}=\frac{\alpha e^{-\alpha t}}{1-e^{-\alpha}}$ so that
$(\dot{m}+\kappa m)\dot{m}$ $=\frac{-\frac{2\kappa}{n+1}e^{-\alpha t}+\kappa}{(1-e^{-\alpha})^{2}}\alpha e^{-\alpha t}$ (A.16)
$=\frac{-\frac{2\kappa\alpha}{n+1}e^{-2\alpha t}}{(1-e^{-\alpha})^{2}}+\frac{\kappa\alpha e^{-\alpha t}}{(1-e^{-\alpha})^{2}}$ (A.17)
Therefore the indefinite integral is given by
$\int(\dot{m}+\kappa\,m)\dot{m}\,\mathrm{dt}=\frac{\frac{\kappa}{n+1}e^{-2\alpha t}-\kappa e^{-\alpha t}}{(1-e^{-\alpha})^{2}}$ (A.18)
and the definite integral:
$\begin{split}\int_{0}^{1}(\dot{m}+\kappa\,m)\dot{m}\,\mathrm{dt}&=\frac{\frac{\kappa}{n+1}e^{-2\alpha t}-\kappa e^{-\alpha t}}{(1-e^{-\alpha})^{2}}\,\Bigg{|}_{0}^{1}\\ &=\frac{\frac{\kappa}{n+1}e^{-2\alpha}-\kappa e^{-\alpha}}{(1-e^{-\alpha})^{2}}-\frac{\frac{\kappa}{n+1}-\kappa}{(1-e^{-\alpha})^{2}}\\ &=\frac{\kappa\left(\frac{1}{n+1}e^{-2\alpha}-e^{-\alpha}-\frac{1}{n+1}+1\right)}{(1-e^{-\alpha})^{2}}\end{split}$ (A.19)
thus
$\frac{\kappa\left(\frac{1}{n+1}(e^{-2\alpha}-1)+(1-e^{-\alpha}\right)}{(1-e^{-\alpha})^{2}}$ (A.20)
However, since $e^{-2\alpha}-1=(e^{-\alpha}-1)(e^{-\alpha}+1)$ the above expression simplifies to
$\frac{\frac{\kappa}{n+1}(e^{-\alpha}+1)+\kappa}{1-e^{-\alpha}}$ (A.21)
and the proposition is proved. $\blacksquare$
A.4 Proof of equilibrium cost of trading
In this section we compute the cost of trading to an a specific trader in an equilibrium. We first note that if $\lambda_{i}$, $a_{i}$ are equilibrium strategies with the market strategy $\sum_{i}\lambda_{i}a_{i}$ we have:
$\text{Cost}(\lambda_{i}a_{i},m)=\lambda_{i}\int_{0}^{1}(\dot{m}+\kappa m)\dot{a}_{i}\,\mathrm{dt}$ (A.22)
We will use equation Eq. 4.8 that says
$\kappa\dot{a}_{i}-\ddot{a}_{i}=\frac{1}{\lambda_{i}}\frac{2\kappa}{n+1}\dot{m}$ (A.23)
as follows. The cost of trading to trader $i$ is the sum of two integrals:
$\text{Cost}(\lambda_{i}a_{i},m)=\lambda_{i}\left(I_{1}+I_{2}\right),\qquad I_{1}=\int_{0}^{1}\dot{m}\,\dot{a}_{i}\,\mathrm{dt},\quad I_{2}=\int_{0}^{1}m\,\dot{a}_{i}\,\mathrm{dt}$ (A.24)
Using integration by parts we compute $I_{1}$:
$I_{1}=m\,\dot{a}_{i}\Big{|}_{0}^{1}-\int_{0}^{1}m\ddot{a}_{i}\,\mathrm{dt}\implies I_{1}+I_{2}=m\,\dot{a}_{i}\Big{|}_{0}^{1}+\int_{0}^{1}m\left(\kappa\dot{a}_{i}-\ddot{a}_{i}\right)$ (A.25)
The integrand in the last part of $I_{1}+I_{2}$ has the term $\kappa\dot{a}_{i}-\ddot{a}_{i}$ and therefore substituting Eq. A.23 we obtain:
$I_{1}+I_{2}=m\,\dot{a}_{i}\Big{|}_{0}^{1}+\frac{1}{\lambda_{i}}\frac{2\kappa}{n+1}\int_{0}^{1}m\dot{m}\,\mathrm{dt}$ (A.26)
and therefore since $\text{Cost}(\lambda_{i}a_{i},m)=\lambda_{i}(I_{1}+I_{2})$ we have
$\text{Cost}(\lambda_{i}a_{i},m)=\lambda_{i}m\cdot\dot{a}_{i}\Big{|}_{0}^{1}+\frac{\kappa}{n+1}m^{2}\Big{|}_{0}^{1}$ (A.27)
But because $m(t)$ is the market strategy we have $m(0)=0$ and $m(1)=1$ so we have
$\text{Cost}(\lambda_{i}a_{i},m)=\lambda_{i}\dot{a}_{i}(1)+\frac{\kappa}{n+1}$ (A.28)
Now examining the expressions for $a_{i}(t)$, $i=1,\dots,n$ in Proposition 5.1 we see that
$\dot{a}_{i}(1)=\kappa B_{i}e^{\kappa}+\alpha D_{i}e^{-\alpha}$ (A.29)
which means
$\text{Cost}(\lambda_{i}a_{i},m)=\lambda_{i}(\kappa B_{i}e^{\kappa}+\alpha D_{i}e^{-\alpha})+\frac{\kappa}{n+1}$ (A.30)
with the constants as in Eq. 5.2:
$B_{i}=\frac{\lambda_{i}n-1}{\lambda_{i}n(e^{\kappa}-1)},\qquad D_{i}=\frac{1}{\lambda_{i}n(1-e^{-\alpha})},\qquad\alpha=\kappa\frac{n-1}{n+1}$ (A.31)
so that the final simplified value is
$\text{Cost}(\lambda_{i}a_{i},m)=\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}$ (A.32)
We immediately see that since $\sum\lambda_{i}=1$, that $\sum(\lambda_{i}n-1)=0$ and thus
Aggregate $\text{Aggregate cost}=\sum\text{Cost}(\lambda_{i}a_{i},m)=\frac{\alpha}{e^{\alpha}-1}+\frac{\kappa n}{n+1}$ (A.33)
And we note that taking the limit as the number of traders grows large yields $\alpha\to\kappa$ and that
$\lim_{n\to\infty}\text{Aggregate cost}=\frac{\kappa}{e^{\kappa}-1}+\kappa$ (A.34)
And this simplifies to
$\frac{\kappa}{e^{\kappa}-1}+\frac{\kappa(e^{\kappa}-1)}{e^{\kappa}-1}=\frac{\kappa e^{\kappa}}{e^{\kappa}-1}=\frac{\kappa}{1-e^{-\kappa}}$ (A.35)
which recovers the same result as Corollary 4.2. $\blacksquare$
A.5 Implementation cost share
Cost share formula can be expressed in terms of a constant term $c_{0}$, a linear term $c_{1}$ that depends on $\lambda_{i}$, and a denominator $d$:
$\begin{split}c_{0}&=-n\left(e^{\kappa}-1\right)\left(e^{\alpha}-1\right)-(n-1)\left(e^{\kappa}-1\right)\\[5.0pt] c_{1}&=-(n+1)n\left(e^{\alpha}-1\right)e^{\kappa}\\[5.0pt] d&=n\left(e^{\kappa}-1\right)\left(-n\left(e^{\alpha}-1\right)-n+1\right)\end{split}$ (A.36)
Factoring out $\left(e^{\kappa}-1\right)$, we get:
$\begin{split}c_{0}&=\left(e^{\kappa}-1\right)\left[-n\left(e^{\alpha}-1\right)-(n-1)\right]=\left(e^{\kappa}-1\right)\left(1-ne^{\alpha}\right)\\[5.0pt] c_{1}&=-(n+1)\cdot(\lambda_{i}\cdot n-1)\cdot(e^{\alpha}-1)\cdot e^{\kappa}\\ d&=n\left(e^{\kappa}-1\right)\left[-n\left(e^{\alpha}-1\right)-n+1\right]=n\left(e^{\kappa}-1\right)\left(-ne^{\alpha}+1\right).\end{split}$ (A.37)
Dividing $c_{0}$ by $d$, we obtain:
$\frac{c_{0}}{d}=\frac{\left(e^{\kappa}-1\right)\left(-ne^{\alpha}+1\right)}{n\left(e^{\kappa}-1\right)\left(-ne^{\alpha}+1\right)}=\frac{1}{n}.$ (A.38)
Dividing $c_{1}$ by $d$, we get:
$\frac{c_{1}}{d}=\frac{-(n+1)\cdot(\lambda_{i}\cdot n-1)\cdot(e^{\alpha}-1)\cdot e^{\kappa}}{n\left(e^{\kappa}-1\right)\left(-ne^{\alpha}+1\right)}$ (A.39)
and the cost share formula becomes
$\frac{c_{0}}{d}+\frac{c_{1}}{d}=\frac{1}{n}+\frac{c_{1}}{d}$ (A.40)
Given the expression:
$\frac{c_{1}}{d}=\frac{-(n+1)(\lambda_{i}n-1)(e^{\alpha}-1)e^{\kappa}}{n(e^{\kappa}-1)(1-ne^{\alpha})}$ (A.41)
We want to express this as $A+B\lambda_{i}$. After separating and simplifying, we get:
$\frac{c_{1}}{d}=A+B\lambda_{i}$ (A.42)
where:
$A=\frac{(n+1)(e^{\alpha}-1)e^{\kappa}}{n(e^{\kappa}-1)(1-ne^{\alpha})}$ (A.43)
and
$B=\frac{-(n+1)(e^{\alpha}-1)e^{\kappa}}{(e^{\kappa}-1)(1-ne^{\alpha})}$ (A.44)
and therefore the cost share formula becomes
$\frac{1}{n}+\frac{(n+1)(e^{\alpha}-1)}{n(1-e^{-\kappa})(1-ne^{\alpha})}-\frac{(n+1)(e^{\alpha}-1)}{(1-e^{-\kappa})(1-ne^{\alpha})}\lambda_{i}$ (A.45)
and this proves the cost share formula. $\blacksquare$